# Contrastive Learning on Multimodal Analysis of Electronic Health Records

### v1

@@ Abstract
Electronic health record (EHR) systems contain a wealth of multimodal clinical data including structured data like clinical codes and unstructured data such as clinical notes. However, many existing EHR-focused studies has traditionally either concentrated on an individual modality or merged different modalities in a rather rudimentary fashion. This approach often results in the perception of structured and unstructured data as separate entities, neglecting the inherent synergy between them. Specifically, the two important modalities contain clinically relevant, inextricably linked and complementary health information. A more complete picture of a patient’s medical history is captured by the joint analysis of the two modalities of data. Despite the great success of multimodal contrastive learning on vision-language, its potential remains under-explored in the realm of multimodal EHR, particularly in terms of its theoretical understanding. To accommodate the statistical analysis of multimodal EHR data, in this paper, we propose a novel multimodal feature embedding generative model and design a multimodal contrastive loss to obtain the multimodal EHR feature representation. Our theoretical analysis demonstrates the effectiveness of multimodal learning compared to single-modality learning and connects the solution of the loss function to the singular value decomposition of a pointwise mutual information matrix. This connection paves the way for a privacy-preserving algorithm tailored for multimodal EHR feature representation learning. Simulation studies show that the proposed algorithm performs well under a variety of configurations. We further validate the clinical utility of the proposed algorithm in real-world EHR data.
@@ 1 Introduction
The growing accessibility of Electronic Health Record (EHR) data presents numerous opportunities for clinical research, ranging from patient profiling (Halpern et al.,, 2016) to predicting medical events (Choi et al.,, 2017). However, the complexity increases with the multimodal nature of EHR data, which encompasses diverse clinical data from patient demographics and genetic information to unstructured textual data like clinical notes, and structured data such as diagnostic and procedure codes, medication orders, and lab results.
@@
A key challenge in EHR-focused research lies in effectively merging these different data types and ensuring that their clinical aspects are meaningfully represented. Research has shown the benefits of integrating structured and unstructured data for tasks like automated clinical code assignment (Scheurwegs et al.,, 2016), managing chronic diseases (Sheikhalishahi et al.,, 2019), and pharmacovigilance (Stang et al.,, 2010). While these different modalities serve as complementary data sources, there is significant overlap and correlation among these data (Qiao et al.,, 2019). Joint representation of both structured and narrative data into a more manageable low-dimensional space where similar features are grouped closely can significantly improve the utility of both data types. This representation learning technique has gained popularity for its ability to capture and represent the intricate relationships among various EHR features.
@@
Despite the extensive research on EHR feature representation, most existing studies have primarily focused on either structured (Choi et al., 2016a, ; Kartchner et al.,, 2017; Hong et al.,, 2021; Zhou et al.,, 2022) or unstructured data modalities (De Vine et al.,, 2014; Choi et al., 2016b, ; Beam et al.,, 2019; Alsentzer et al.,, 2019; Huang et al.,, 2020; Lehman and Johnson,, 2023). For instance, Alsentzer et al., (2019) adapted the BERT model (Devlin et al.,, 2019) to the clinical domain by training on the MIMIC-III clinical notes (Johnson et al.,, 2016). It was extremely time/resource-consuming to train such a black-box model. De Vine et al., (2014) aligned free-text features with the Unified Medical Language System (UMLS)’s common concept unique identifier (CUI) space (McInnes et al.,, 2007). They then utilized the skip-gram algorithm (Mikolov et al.,, 2013) on concatenated concept documents to learn the CUI embeddings. Similarly, Choi et al., 2016b and Beam et al., (2019) employed singular value decomposition (SVD) on a pointwise mutual information (PMI) matrix derived from CUI co-occurrences in unstructured text to generate CUI embeddings. This approach was also adopted by Hong et al., (2021) for code embeddings. This SVD-PMI algorithm not only ensures scalability but also upholds data privacy through the use of aggregate co-occurrence data, offering a scalable variant of the skip-gram algorithm (Levy and Goldberg,, 2014). Its interpretability is further highlighted in the dynamic log-linear topic model developed by Arora et al., (2016).
Recent studies have emerged on leveraging multimodal EHR features for enhanced predictive modeling, as demonstrated by the work of Khadanga et al., (2019); Zhang et al., (2020); Bardak and Tan, (2021); Gan et al., (2023). Specifically, Gan et al., (2023) enhanced code and CUI embeddings by employing the SVD-PMI algorithm, facilitating the integration of structured and unstructured data. Despite these advancements, methods relying on straightforward data merging may not fully account for the intricate interactions within multimodal data, potentially leading to biases. This issue will be further examined in our theoretical (Section 3) and numerical analyses (Sections 4 and 5).
Addressing these limitations, Liu et al., (2022) introduced a multimodal pre-trained language model incorporating a cross-attention mechanism to enrich EHR representations across structured and unstructured data domains. Additionally, the domain has seen progress in multimodal contrastive learning strategies, as evidenced by research from Li and Gao, (2022); Yin et al., (2023); Wang et al., (2023). These strategies, drawing inspiration from successful vision-language models like the Contrastive Language-Image Pre-Training (CLIP) by Radford et al., (2021), aim to train unified representations of diverse data modalities. However, these approaches, grounded in deep neural networks, face challenges due to their “black-box” nature, including a lack of theoretical underpinning, computational complexity, and concerns over privacy since they necessitate access to individual patient data. These factors contribute to their limited applicability in the context of EHR data.
@@
While some theoretical analyses of multimodal learning exist, their applicability to EHR data has been limited. Groundbreaking studies like Huang et al., (2021) have illustrated the benefits of multimodal learning, showing that learning across multiple modalities can reduce population risk compared to single-modality methods. Recently, Deng et al., (2023) theoretically proved the zero-shot transfer ability of CLIP. Furthermore, Nakada et al., (2023) explored multimodal contrastive learning’s performance under a spiked covariance model. However, these studies do not directly apply to the unique discrete feature structure of EHR data, leaving an unaddressed theoretical gap in understanding multimodal contrastive learning’s application in healthcare. Bridging this gap is vital, as it lays a solid foundation for multimodal contrastive learning’s development and implementation in healthcare, maximizing its potential to improve patient care and drive medical research forward.
@@
To overcome these limitations, we introduce the Contrastive Learning Algorithm for Integrated Multimodal Electronic health records (CLAIME). Our findings confirm that CLAIME is not only an effective tool for deriving multimodal EHR feature representations but also a method that respects privacy by requiring only aggregated data. Additionally, we propose a novel multimodal feature embedding generative model (2.1) in Section 2.1, designed to enhance statistical analysis of multimodal EHR data. This model is notable for its interpretability and accurate portrayal of EHR data generation. It distinguishes itself from earlier word vector generative models (Arora et al.,, 2016, 2018; Lu et al.,, 2023; Xu et al.,, 2022) by (1) enabling the integration of multimodal EHR features, and (2) allowing patient heterogeneity by incorporating error terms specific to patients, thus increasing the model’s robustness. Within this generative framework, we validate the consistency of the CLAIME algorithm and clarify the relationship between multimodal feature embeddings and a multimodal PMI matrix. The proposed algorithm is also privacy-preserving since it only requires summary-level data, opening doors for collaboration across multiple institutions. Our research also fills a theoretical void in the analysis of multimodal contrastive learning for EHR data.
The rest of the paper is structured as follows. Section 2 elaborates on the proposed method. Section 3 presents the theoretical properties of our algorithm. In Section 4, simulation studies demonstrate the algorithm’s effectiveness under various configurations. Section 5 further validates our algorithm’s clinical applicability in EHR studies. Finally, Section 6 includes discussions.
@@ 2 Method
2.1 Notation
For any matrix $\mathbf{A}$, let $\|\mathbf{A}\|$, $\|\mathbf{A}\|_{\max}$ and $\|\mathbf{A}\|_{\mathrm{F}}$ be its operator norm, entrywise maximum norm and Frobenius norm, respectively. We define $\mathbf{P}_{p}(\mathbf{A})$ as the top-$p$ right singular vectors of $\mathbf{A}$. When the right singular vectors are not unique, we choose arbitrary singular vectors. Denote $s_{j}(\mathbf{A})$ as the $j$-th largest singular value of $A$. Let $\mathcal{O}_{d,p}$ ($d\geq p$) denote a set of $d\times p$ orthonormal matrices. For two sequences of positive numbers $\{a_{k}\}$ and $\{b_{k}\}$, we write $a_{k}\lesssim b_{k}$ or $a_{k}=O(b_{k})$ or $b_{k}\gtrsim a_{k}$ or $b_{k}=\Omega(a_{k})$ if there exists some constant $C>0$ such that $a_{k}\leq Cb_{k}$ for all $k$. We denote $a_{k}\ll b_{k}$ or $a_{k}=o(b_{k})$ if $\lim_{k\rightarrow\infty}a_{k}/b_{k}=0$. For any positive integer $I$, let $[I]=\{1,2,\cdots,I\}$. We write $a\vee b$ and $a\wedge b$ to denote $\max(a,b)$ and $\min(a,b)$, respectively. We denote by $\bm{e}_{j}$ the $j$-th unit vector in $\mathbb{R}^{d}$ where its $j$-th element is one, and all other elements are zero.
@@ 2.2 Model Assumptions
Assume that the collections of codes from the structured modality and CUIs from the unstructured modality are denoted by ${\cal W}^{(1)}:=[d_{1}]$ and ${\cal W}^{(2)}:=[d]\setminus[d_{1}]$ respectively, where $d=d_{1}+d_{2}$, with $d$ representing the total number of unique features across both modalities. Suppose that we have $n$ independent patients. For each patient $i$, the observed codes and CUIs are denoted as $\{w^{(1)}_{i,t}\}_{t\in[T_{i}^{(1)}]}$ and $\{w^{(2)}_{i,t}\}_{t\in[T_{i}^{(2)}]}$, where $w^{(1)}_{i,t}\in{\cal W}^{(1)}$ and $w^{(2)}_{i,t}\in{\cal W}^{(2)}$. The sizes of these sets are given by $T_{i}^{(1)}\geq 2$ and $T_{i}^{(2)}\geq 2$, respectively. We model the probability of observing specific codes and CUIs for the $i$-th patient based on their embeddings $\mathbf{v}_{w}^{\star}\in\mathbb{R}^{p}$ as follows:
$\mathbb{P}(w^{(1)}_{i,t}=w|\bm{c}_{i},{\bm{\epsilon}}_{i}^{(1)})$ $=\frac{{\exp}(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w})}{\sum_{w^{\prime}\in{\cal W}^{(1)}}{\exp}(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}})},\ \ w\in{\cal W}^{(1)},t\in[T_{i}^{(1)}],\text{ and }$ (2.1)
@@
$\mathbb{P}(w^{(2)}_{i,t}=w|\bm{c}_{i},{\bm{\epsilon}}_{i}^{(2)})$ $=\frac{{\exp}(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w})}{\sum_{w^{\prime}\in{\cal W}^{(2)}}{\exp}(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}})},\ \ w\in{\cal W}^{(2)},t\in[T_{i}^{(2)}]\,.$
Here $\bm{c}_{i}\sim N(\bm{0},\mathbf{I}_{p})$ represents a latent vector specific to patient $i$, reflecting their clinical state. The error terms, ${\bm{\epsilon}}^{(1)}_{i}=(\epsilon_{i,1},\dots,\epsilon_{i,d_{1}})^{\top}\sim N(\bm{0},{\bm{\Sigma}}_{1})$ and ${\bm{\epsilon}}^{(2)}_{i}=(\epsilon_{i,d_{1}+1},\dots,\epsilon_{i,d})^{\top}\sim N(\bm{0},{\bm{\Sigma}}_{2})$, account for patient-specific variations and address the variability not captured by the term $\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle$, where ${\bm{\Sigma}}_{1}$ and ${\bm{\Sigma}}_{2}$ are some unknown positive semi-definite matrices.
2.3 CLAIME Algorithm

### v2

@@ Abstract
Electronic health record (EHR) systems capture a wealth of multimodal clinical data, encompassing both structured clinical codes and unstructured clinical notes. Yet, many EHR-focused studies have traditionally examined these modalities in isolation or combined them using simplistic methods. Such approaches reinforce the view of structured and unstructured data as separate, overlooking the intrinsic synergy between them. In reality, these modalities are deeply interconnected, each containing clinically relevant and complementary information that, when integrated effectively, can provide a more comprehensive understanding of patient health. Integrating these modalities enables a richer, more complete representation of a patient’s medical history, offering deeper insights into their health. Despite the great success of multimodal contrastive learning on vision-language, its potential remains under-explored in the realm of multimodal EHR, particularly in terms of its theoretical understanding. To support statistical analysis of multimodal EHR data, we propose a novel multimodal feature embedding generative model and design a multimodal contrastive loss to obtain the multimodal EHR feature representation. Our theoretical analysis demonstrates the effectiveness of multimodal learning compared to single-modality learning and connects the solution of the loss function to the singular value decomposition of a pointwise mutual information matrix. This connection paves the way for a privacy-preserving algorithm tailored for multimodal EHR feature representation learning. Simulation studies show that the proposed algorithm performs well under a variety of configurations. We further validate the clinical utility of the proposed algorithm in real-world EHR data.
@@ 1 Introduction
The growing accessibility of Electronic Health Record (EHR) data presents numerous opportunities for clinical research, ranging from patient profiling (Halpern et al., , 2016) to predicting medical events (Choi et al., , 2017). However, the complexity increases with the multimodal nature of EHR data, which encompasses diverse data from patient demographics and genetic information to unstructured textual data like clinical notes, as well as structured data such as diagnostic and procedure codes, medication orders, and lab results.
@@
A key challenge in EHR-focused research lies in effectively merging these different data types and ensuring that their clinical aspects are meaningfully and coherently represented. Research has shown the benefits of integrating structured and unstructured data for tasks like automated clinical code assignment (Scheurwegs et al., , 2016), managing chronic diseases (Sheikhalishahi et al., , 2019), and pharmacovigilance (Stang et al., , 2010). While these different modalities serve as complementary data sources, there is significant overlap and correlation among these modalities (Qiao et al., , 2019). Joint representation of both structured and narrative data into a more manageable low-dimensional space where similar features are grouped closely can significantly improve the utility of both data types. This representation learning technique has gained popularity for its ability to capture and represent the intricate relationships among various EHR features.
@@
Despite extensive research on EHR feature representation, most studies have focused on datasets with a single modality, including structured (Choi et al., 2016a, ; Kartchner et al., , 2017; Hong et al., , 2021; Zhou et al., , 2022) and unstructured data (De Vine et al., , 2014; Choi et al., 2016b, ; Beam et al., , 2019; Alsentzer et al., , 2019; Huang et al., , 2020; Lehman and Johnson, , 2023). For example, Alsentzer et al., (2019) adapted BERT (Devlin et al., , 2019) to the clinical domain by pretraining on MIMIC-III notes (Johnson et al., , 2016), a process that was highly resource-intensive. De Vine et al., (2014) aligned free-text with the UMLS Concept Unique Identifier (CUI) space (McInnes et al., , 2007) and applied skip-gram (Mikolov et al., , 2013) to learn embeddings, while Choi et al., 2016b , Beam et al., (2019); Hong et al., (2021); Levy and Goldberg, (2014); Arora et al., (2016) used singular value decomposition (SVD) on pointwise mutual information (PMI) matrices.
Building on these unimodal advances, recent work has explored multimodal EHR representations that integrate both structured and unstructured information (Khadanga et al., , 2019; Zhang et al., , 2020; Bardak and Tan, , 2021; Gan et al., , 2025). Joint representation is preferable to treating all data as a single modality because, although structured codes and unstructured text encode overlapping clinical concepts, they possess fundamentally different statistical forms. Ignoring these cross-modality differences can obscure modality-specific information, discard complementary signal, and amplify systematic bias (see further discussion in Section 3.2). By contrast, multimodal integration can exploit the complementary strengths of distinct modalities, yielding representations that are not only richer but also less biased than those obtained from unimodal analyses. For instance, information that is noisy or missing in one modality may be corroborated by the other, thereby stabilizing estimation and enhancing generalization. In this paper, we develop a theoretical framework (Section 3) and provide empirical evidence (Sections 4 and 5) to demonstrate that multimodal analysis addresses these challenges and produces more reliable and interpretable representations of EHR data.
To address these limitations, Liu et al., (2022) proposed a multimodal pre-trained language model that incorporates a cross-attention mechanism to enrich EHR representations across structured and unstructured domains. More recently, multimodal contrastive learning strategies have also been explored (Li and Gao, , 2022; Yin et al., , 2023; Wang et al., , 2023). Inspired by the success of vision–language models such as CLIP (Radford et al., , 2021), these approaches aim to learn unified representations of heterogeneous data modalities. From a statistical perspective, contrastive learning can be interpreted through the lens of noise-contrastive estimation (Gutmann and Hyvärinen, , 2010), in which a parametric score function distinguishes positive’ pairs drawn from the joint distribution from negative’ pairs drawn from the product of marginals. The popular InfoNCE loss (Oord et al., , 2018) maximizes a lower bound on mutual information and is closely related to low-rank approximations of the PMI matrix (Levy and Goldberg, , 2014). Despite their promise, these methods remain limited by their reliance on deep neural networks, which are often criticized for their “black-box” nature. Practical challenges include the absence of rigorous theoretical guarantees, high computational cost, and privacy concerns due to the need for patient-level data. Together, these issues constrain the applicability of existing multimodal contrastive learning frameworks in the context of EHR data.
@@
While some theoretical analyses of multimodal learning exist, their applicability to sparse discrete EHR data has been limited. Groundbreaking studies like Huang et al., (2021) have illustrated the benefits of multimodal learning, showing that learning across multiple modalities can reduce population risk compared to single-modality methods. Recently, Deng et al., (2023) theoretically proved the zero-shot transfer ability of CLIP. Furthermore, Nakada et al., (2023) explored multimodal contrastive learning’s performance under a spiked covariance model. However, these studies do not directly apply to the unique discrete feature structure of EHR data, leaving an unaddressed theoretical gap in understanding multimodal contrastive learning’s application in healthcare. Bridging this gap is vital, as it lays a foundation for multimodal contrastive learning’s development and implementation in healthcare, maximizing its potential to improve care and drive medical research forward.
Complementary lines of research have attempted to address related challenges from different angles. For instance, exponential family principal component analysis (PCA) (Collins et al., , 2001) and nonlinear PCA (Hinton and Salakhutdinov, , 2006) extend classical PCA to non-Gaussian and nonlinear settings, while kernel PCA (KPCA) (Gupta et al., , 2019) provides additional flexibility. Yet, handling multimodal, sparse, and privacy-sensitive data remains an open challenge. In parallel, the statistical literature on integrative dimension reduction, such as JIVE (Lock et al., , 2013), AJIVE (Feng et al., , 2018), and other multi-view factor models (Zhou et al., , 2015; Yang and Michailidis, , 2016; Yi et al., , 2023, e.g.,), provides tools for modeling shared and modality-specific structures. Our framework shares this goal but is tailored to discrete EHR features and prioritizes privacy-preserving learning via co-occurrence data.
@@
To overcome these limitations, we introduce the Contrastive Learning Algorithm for Integrated Multimodal Electronic health records (CLAIME). Our findings confirm that CLAIME is not only effective for deriving multimodal EHR feature representations but also respects privacy by requiring only aggregated data. Additionally, we propose a novel multimodal feature embedding generative model (2.1) in Section 2.1, designed to enhance statistical analysis of multimodal EHR data. This model is notable for its interpretability and accurate portrayal of EHR data generation. It distinguishes itself from earlier word vector generative models (Arora et al., , 2016, 2018; Lu et al., , 2023; Xu et al., , 2023) by (1) enabling the integration of multimodal EHR features, and (2) allowing patient heterogeneity by incorporating error terms specific to patients, thus increasing the model’s robustness. Within this generative framework, we validate the consistency of the CLAIME algorithm and clarify the relationship between multimodal feature embeddings and a multimodal PMI matrix. The proposed algorithm is also privacy-preserving since it only requires summary-level data, opening doors for collaboration across multiple institutions. Our research also fills a theoretical void in the analysis of multimodal contrastive learning for EHR data.
The rest of the paper is organized as follows. Section 2 introduces the proposed method, and Section 3 presents its theoretical properties. Sections 4 and 5 provide simulation and real-world evaluations, respectively. Section 6 concludes with a discussion.
@@ 2 Method
We begin by defining the notation. For any matrix $\mathbf{A}$, let $\|\mathbf{A}\|$, $\|\mathbf{A}\|_{\max}$, and $\|\mathbf{A}\|_{\mathrm{F}}$ denote its operator norm, entrywise maximum norm, and Frobenius norm, respectively. Let $\mathbf{P}_{p}(\mathbf{A})$ be the top-$p$ right singular vectors of $\mathbf{A}$, chosen arbitrarily if non-unique, and let $s_{j}(\mathbf{A})$ be its $j$-th largest singular value. Define $\mathcal{O}_{d,p}$ ($d\geq p$) as the set of $d\times p$ orthonormal matrices. For sequences $\{a_{n}\},\{b_{n}\}>0$, write $a_{n}\lesssim b_{n}$, $a_{n}=O(b_{n})$, $b_{n}\gtrsim a_{n}$, or $b_{n}=\Omega(a_{n})$ if $a_{n}\leq Cb_{n}$ for some $C>0$ and all $n$; write $a_{n}\ll b_{n}$ or $a_{n}=o(b_{n})$ if $a_{n}/b_{n}\to 0$. Let $I]=\{1,2,\dots,I\}$ for any positive integer $I$, and write $a\vee b$ and $a\wedge b$ for $\max(a,b)$ and $\min(a,b)$, respectively. Denote by $\bm{e}_{j}$ the $j$-th unit vector in $\mathbb{R}^{d}$.
@@ 2.1 Model Assumptions
[…]
Figure 1: Illustration of multimodal EHR generation: A shared latent state $\bm{c}_{i}$ generates both structured codes (top) and unstructured CUIs (bottom), motivating (2.1).
Our approach is motivated by real-world EHRs, where structured codes (e.g., diagnoses, medications) and unstructured concepts extracted from clinical notes (e.g., CUIs) offer complementary views of a patient’s health (Figure 1). During a hospital visit, patient $i$ generates two types of data: structured codes and CUIs, which are typically collected by different subsystems and subject to distinct noise processes. For example, structured codes often reflect billing practices, while CUIs are derived from free-text notes via Named Entity Recognition (NER) tools. This data processing pipeline aligns with standard practice for mapping text to UMLS concepts (Choi et al., 2016b, ; Beam et al., , 2019). We primarily focus on the two-modality case but also discuss extensions to more general multi-modality settings involving three or more modalities below.
Accordingly, for each patient $i$, we observe two sets of features: $\{w^{(1)}_{i,t}\}_{t=1}^{T_{i}^{(1)}}$ for structured codes and $\{w^{(2)}_{i,t}\}_{t=1}^{T_{i}^{(2)}}$ for CUIs, where $w^{(1)}_{i,t}\in{\cal W}^{(1)}$ and $w^{(2)}_{i,t}\in{\cal W}^{(2)}$. Let ${\cal W}^{(1)}:=[d_{1}]$ and ${\cal W}^{(2)}:=[d]\setminus[d_{1}]$, where $d=d_{1}+d_{2}$ denotes the total number of distinct features across both modalities. To model the generative process of these features, we adopt the following latent variable model, which captures the shared structure between modalities while allowing for modality-specific noise:
@@
$\mathbb{P}(w^{(1)}_{i,t}=w\mid\bm{c}_{i},{\bm{\epsilon}}_{i}^{(1)})$ $=\frac{\exp(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w}^{(1)})}{\sum_{w^{\prime}\in{\cal W}^{(1)}}\exp(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}}^{(1)})},\quad w\in{\cal W}^{(1)},\ t\in[T_{i}^{(1)}],$ (2.1)
$\mathbb{P}(w^{(2)}_{i,t}=w\mid\bm{c}_{i},{\bm{\epsilon}}_{i}^{(2)})$ $=\frac{\exp(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w}^{(2)})}{\sum_{w^{\prime}\in{\cal W}^{(2)}}\exp(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}}^{(2)})},\quad w\in{\cal W}^{(2)},\ t\in[T_{i}^{(2)}].$
Here, $\bm{c}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{p})$ represents the latent clinical state of patient $i$, and $\mathbf{v}_{w}^{\star}\in\mathbb{R}^{p}$ is the ground-truth embedding of feature $w$. The modality-specific noise vectors ${\bm{\epsilon}}^{(M)}_{i}=(\epsilon_{i,w}^{(M)})_{w\in{\cal W}^{(M)}}\sim\mathcal{N}(\mathbf{0},{\bm{\Sigma}}_{M})$, for $M\in\{1,2\}$, capture variability not explained by the latent state, such as documentation inconsistencies or NLP extraction errors.
This formulation reflects our core assumption: while structured and unstructured features are conditionally generated from a shared latent representation $\bm{c}_{i}$, they are distorted independently by modality-specific noise. The model thus supports: (1) integration of heterogeneous EHR data sources, (2) robustness to differing noise structures across modalities, and (3) inference of informative embeddings from only summary-level statistics.
Remark 2.1.
The assumption $\bm{c}_{i}\sim N(\mathbf{0},\mathbf{I}_{p})$ is used to support theoretical analysis, not as a requirement for the CLAIME algorithm itself. Our method does not rely on estimating individual $\bm{c}_{i}$ vectors. Instead, the Gaussian assumption enables analytical tractability in characterizing the expected co-occurrence patterns of EHR features, which in turn justify the use of SVD-based estimators in later sections. Similar analytical approximations have been derived under alternative latent structures, such as unit sphere random walks (Arora et al., , 2016, 2017) or scaled-Gaussian priors (Xu et al., , 2023). Finally, due to the rotational invariance of our model, the identity covariance assumption on $\bm{c}_{i}$ can be relaxed without affecting the core methodology.
2.2 The CLAIME Algorithm